Calculate the volume of hydrogen gas produced at STP when 30g of Zn is added to dilute hydrochloric acid at 33⁰c and 775mmHg pressure (H=1 , Zn=65 , cl =35.5) molar mass of gas at STP= 22.4 dm³

1
Expert's answer
2022-04-15T00:42:03-0400

Zn + 2HCl = ZnCl2 + H2

n(Zn) = 30 g / 65g/mol = 0.4615 mol

1 mol Zn - 1 mol H2

0.4615 mol Zn - x mol H2

n(H2) = x = 0.4615 mol.

V(H2) = n(H2) × Vm = 0.4615 mol × 22.4 L/mol = 10.34 L